verify that each given function is a solution of the

differential equation.

6.′′\:+\:2y′\:−\:3y\:=\:0;\:y\left(t\right)\:=\:e^{−3t}


1
Expert's answer
2021-09-02T23:14:06-0400

"y′′\:+\:2y′\:−\:3y\:=\:0;\:y\left(t\right)\:=\:e^{−3t}"

Replacing "y'" by "D" , "y''" by "D^2" , we get:

"D^2+2D-3=0\\
\Rightarrow D^2+3D-D-3=0\\
\Rightarrow D(D+3)-1(D+3)=0\\
\Rightarrow (D-1)(D+3)=0\\
\therefore D=1,-3\\"

"\frac{1}{y}\frac{dy}{dx}=1,-3\\
\Rightarrow \frac{dy}{y}=dx; \frac{dy}{y}=-3dx\\"

Integrating both sides, we get:

"ln(y)=x+c,-3x+c"

"\therefore y=ke^x,ke^{-3x}"


Now verifying the solution:

When "y=ke^x," we get:

"\frac{dy}{dx}=ke^x,\frac{d^2y}{dx^2}=ke^x"

Now, substituting, these in the given differential equation, we get:

"y''+2y'-3y=0\\
\Rightarrow ke^x+2ke^x-3ke^x=0 \\
\Rightarrow 0=0"

Hence Verified.


When "y=ke^{-3x}" , we get:

"\frac{dy}{dx}=-3ke^{-3x},\frac{d^2y}{dx^2}=9ke^{-3x}"

Now, substituting, these in the given differential equation, we get:

"y''+2y'-3y=0\\
\Rightarrow 9ke^{-3x}-6ke^{-3x}-3ke^{-3x}=0 \\
\Rightarrow 0=0"

Hence Verified.


Hence, solutions are obtained and verified.